Solve by substitution:
3x-y=5
-6x +2y=1
I dont know how to start this. Can you help me solve this?

Answer by stanbon(75887)

Solve by substitution:
3x-y=5
-6x +2y=1
===============
Solve the 1st equation for "y":
y = 3x-5
----
Substitute that value for "y" into the 2nd equation and solve for "x":
---
-6x + 2(3x-5) = 1
-6x+6x-10 = 2
-10 = 2
---
That is a contradiction and tells you there
is no solution for the system of equations.
---
Note: If you would graph the two equations
you would see the graphs are parallel; they
have no point in common.
Cheers,
Stan H.